# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ 4.3 Concrete Choices of $L$, $m$ and ${\sf pos}$
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed to obtain $L$ and $m$ is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that $\beta,\gamma>2^{\kappa}$. This ensures security against (fast) algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (21), (22) and (23) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. We note, however, that generating $2^{\kappa}$ keystream bits from a single key and IV pair is meaningless from a practical point of view. Instead, we set $B=64$, i.e. from a single key and IV pair at most $2^{64}$ keystream bits are to be generated. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
$\kappa$ $L$ $m$ $12m+$ 1 deg ${\sf LB}$ ${\sf AI}$ ${\sf FAI}$
@@
    The number of variables of the filtering function $f_{2m+1}$ is $12m+$ 1 (and not $m$).
@@
    For each entry in the table, our calculation shows that $\beta\gg\gamma$. For example, for $\mathcal{S}(257,59)$, we obtained $\beta=2^{364.38}$ while $\gamma=2^{130.12}$. This is not surprising, since the fast algebraic attack is known to be much more faster than the basic algebraic attack.
@@
    The feedback polynomial for the LFSR has to be a primitive polynomial of degree $L$. The above analysis of correlation attacks shows that for the obtained values of $L$ and $m$, it does not matter whether the polynomial is sparse, or whether it has a sparse multiple. So from an efficiency point of view, one may choose a low weight primitive polynomial. Examples of primitive polynomials for the values of $L$ in Table 1 are shown in Table 2.
@@
Table 2: Examples of primitive polynomials for the values of $L$ shown in Table 1.
@@
Given the values of $L$ and $m$ (and also $\kappa$), we applied the procedure for choosing the tap positions encoded by ${\sf pos}$ as outlined in Section 4.1. The leftmost $\kappa$ bits of ${\sf pos}$ is the string ${\sf posX}$. The next $\kappa-1$ bits of ${\sf pos}$ (i.e. the positions ${\sf pos}[L-\kappa-1,\ldots,L-2\kappa+1]$) encode the tap positions for the variables in $\mathbf{Y}$. We define ${\sf posY}$ to be the $\kappa$-bit string formed by appending a 0 to the $(\kappa-1)$-bit segment ${\sf pos}[L-\kappa-1,\ldots,L-2\kappa+1]$ of ${\sf pos}$. So both ${\sf posX}$ and ${\sf posY}$ are $\kappa$-bit strings and are uniquely defined from ${\sf pos}$. Further, given ${\sf posX}$ and ${\sf posY}$ it is possible to uniquely construct ${\sf pos}$ (by concatenating ${\sf posX}$ and ${\sf posY}$, appending $L-2m$ zeros, and setting the bit in position $L-2\kappa$ to be 1, to encode the tap position for $W$). So providing ${\sf posX}$ and ${\sf posY}$ is equivalent to providing ${\sf pos}$. In Table 3 we provide the values of ${\sf posX}$ and ${\sf posY}$ encoded as hexadecimal strings. For example, the entry for ${\sf posX}$ corresponding to $\mathcal{S}(128,59)$ is be352$\cdots$ which encodes the binary string 10111110001101010010$\cdots$. In Table 4 we provide the corresponding values of $\nu$ and $\delta$. Recall that $\nu$ is obtained from ${\sf posX}$ as in (17) and $\delta$ is obtained from ${\sf pos}$ as in (18).
@@
$\kappa$ ${\sf posX}$ and ${\sf posY}$
$\mathcal{S}(80,37)$ d569a664f500763506c3
@@
$\mathcal{S}(80,37)$ 16 36
$\mathcal{S}(128,59)$ 26 57
$\mathcal{S}(160,71)$ 30 69
$\mathcal{S}(192,87)$ 39 86
$\mathcal{S}(224,101)$ 45 96
$\mathcal{S}(256,115)$ 51 112
@@ 4.4 Differential Attacks
Choosing the filtering function to protect against certain known classes of attacks does not however, protect against all possible attacks. There are attacks which can succeed even if the filtering function is properly chosen. One such attack on a previous version of the proposal was described in [6]. To describe the attack and the design modification to resist it we need to mention the aspects of the design which were different in the previous proposal. There are three such aspects.
@@
    In the definition of ${\sf MM}_{2m}$, in the earlier version we had chosen $\pi$ to be the identity permutation, whereas in the present version we choose $\pi$ to be the bit reversal permutation.
@@
In other words, due to the choice of $\pi$ as the identity permutation, the selection of tap positions as consecutive positions of the state, and the fact that two successive keystream bits are obtained from a single shift of the LFSR sequence, $m-1$ of the quadratic terms in the inner product cancelled out when $w_{t}$ and $w_{t+1}$ are XORed together. Further, the inputs to the two calls to ${\sf Maj}$ have an overlap of size $m-1$. These features were observed in [6] and exploited to mount a differential attack. The idea of the differential attack is to introduce a difference in the top most bit position of the IV. During the initialisation phase, this difference travels unchanged for $\kappa-m$ steps and then produces a difference in the input to the filtering function. The difference in the output of the filtering function is fed back into the leftmost bit of the LFSR. As the LFSR is further shifted, this difference then travels without any further modification creating a high probability truncated differential. Combined with the simplified form of $w_{t}\oplus w_{t+1}$, this leads to an efficient key recovery attack.
@@
Further, due to our design procedure for the tap positions of corresponding to the variables $X_{1},\ldots,X_{n}$, the overlap of inputs in the calls to ${\sf Maj}$ in the XOR of any two keystream bits is at most $\nu$. From Table 4, we observe that the value of $\nu$ is considerably smaller than $m$. So the complexity of ${\sf Maj}$ is mostly preserved in the XOR of any number of keystream bits.
Further, in the present version, during the initialisation phase we inject the output of the filtering function at multiple positions of the state, the number of positions being controlled by the parameter $\mu$. So any (controlled) difference in any bit position travels at most $\mu$ steps before it is further modified. Since our recommendation is to choose $\mu$ to be about $\sqrt{2\kappa}$, in the full initialisation phase consisting of $2\kappa$ iterations, a difference will be updated about $\sqrt{2\kappa}$ times. This makes it difficult to control a difference through all the initialisation rounds.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
The computation of ${\sf MM}_{2m}$ requires computing $h$ and an inner product of two $m$-bit strings. The computation of $h$ requires the computation of the weight of an $m$-bit string and the computation of a threshold function. We discuss basic strategies for implementing these operations which provide estimates of the number of gates required.
@@
This operation requires $m$ AND gates and $m-1$ XOR gates.
@@
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $12^{r}-r-$ 1 full adders. From this it easily follows that the number of full adders required to compute the weight of an $m$-bit string for arbitrary $m$ is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$. We record this fact in the following result.
@@
Threshold function on the weight of an $m$-bit string.
The weight of a 37-bit string is a 6-bit quantity, say $w_{5}w_{4}w_{3}w_{2}w_{1}w_{0}$. It is required to determine whether the value represented by this string is at least 19. This is computed by the Boolean formula $w_{5}\vee(w_{4}\wedge(w_{3}\vee w_{2}\vee(w_{1}\wedge w_{0})))$, requiring 3[OR]+2[AND] gates. In general, the weight of an $m$-bit string is an $\omega$-bit value, where $\omega=\lceil\log_{2}m\rceil$. To compute the threshold function, $\omega_{1}$ OR and $\omega_{2}$ AND gates are required for some values of $\omega_{1}$ and $\omega_{2}$ satisfying $\omega_{1}+\omega_{2}\leq\omega$. So the number of gates for computing the threshold function on the weight of an $m$-bit string requires a logarithmic (in $m$) number of gates.
@@
Next we consider concrete estimates. For such estimates, we ignore the at most $\lceil\log_{2}m\rceil$ AND and OR gates required for computing the threshold function from the weight, and the gates required to compute the next bit of the LFSR. Further, we also ignore the $\lceil 2\kappa/\mu\rceil$ XOR gates required for the feedback injection from the filtering function during the initialisation phase. This number is not much. For example, for $\kappa=128$, we have $\mu=\lfloor\sqrt{2\kappa}\rfloor=16$ and so 16 XOR gates are required to feedback the output of the filtering function during the initialisation phase. We obtain concrete estimates for the two major components, the LFSR and the filtering function.
@@
If we consider a flip-flop to be 12 units as done for Trivium [9], then the gate count estimate for $\mathcal{S}(163,37)$ is 2395.5. The gate estimate for Trivium obtained in [9] is 3488. Both Trivium and $\mathcal{S}(163,37)$ are targeted at the 80-bit security level, and $\mathcal{S}(163,37)$ is substantially smaller than Trivium. The lower size of $\mathcal{S}(163,37)$ is due to the smaller state size; $\mathcal{S}(163,37)$ uses a 163-bit state, while Trivium uses a 288-bit state.
The gate count estimate for Grain-128a obtained in [1] is 2145.5. Grain-128a is targeted at the 128-bit security level. Comparing to $\mathcal{S}(257,59)$, which is also targeted at the 128-bit security level, we find the gate count estimate of $\mathcal{S}(257,59)$ to be 2771.5. The state sizes of both Grain-128a and $\mathcal{S}(257,59)$ are almost the same. The greater size of $\mathcal{S}(257,59)$ is due to the greater gate size requirement of the filtering function $f_{2m+1}$ in comparison to the gate size requirement of the nonlinear components of Grain-128a. Even though $\mathcal{S}(257,59)$ is larger than Grain-128a, its size of about 2771.5 gates is small enough for $\mathcal{S}(257,59)$ to be considered as a small cipher. Finally we note that $\mathcal{S}(521,114)$ which is targeted at the 256-bit security level requires about 5607.5 gates. We know of no other 256-bit secure stream cipher which has such a small gate count.
We note that it is possible to ramp up security at the cost of a reasonable increase in gate count. For example, from Table 1, at the 128-bit security level, our proposal has $L=257$ and $m=59$, resulting in linear bias equal to $2^{-60}$ and the fast algebraic attack requiring more than $2^{130.12}$ operations. If we increase $m$ from 59 to 63, then the linear bias drops to $2^{-64}$, and the fast algebraic attack now requires more than $2^{135.83}$ operations. The gate estimate for $\mathcal{S}(257,63)$, i.e. the stream cipher with $L=257$ and $m=63$, is 2818.5 gates (the LFSR requires 2056 gates, and $f_{127}$ requires 762.5 gates). So the security increases by about 5 bits at the cost of an increase of only 47 gates in the circuit size.

### v7

@@ 4.3 Concrete Choices of $L$, $m$ and $\sf pos$}
Given $\kappa$, we obtained representative values of $L$ and $m$. The procedure we followed to obtain $L$ and $m$ is to choose the value of $L$ to be the first prime number greater than $2\kappa$, and then for the chosen value of $L$, choose $m$ to be the least integer such that both $\beta$ and $\gamma$ (given by (31) and (32) respectively) are greater than $2^{\kappa}$. This ensures security against (fast) algebraic attacks considered above. Next, using $\kappa$ and the corresponding values of $L$ and $m$, we computed the maximum value of $B$ such that $B\leq\kappa$, $\alpha>2^{B}$, and (28), (29) and (30) hold. In each of the cases that we considered, it turns out that this maximum value of $B$ is in fact $\kappa$. So for the chosen values of $L$ and $m$, the corresponding $\mathcal{S}(L,m)$ ensures $\kappa$-bit security against the above analysed correlation attacks even when the adversary has access to $2^{\kappa}$ keystream bits. Note that we allow at most $2^{64}$ keystream bits to be generated from a single key and IV pair. Since $64<\kappa$, it follows that the concrete proposals are indeed secure against that fast correlation attacks considered above. Table 1 shows the values of $\kappa$, $L$, $m$ and other parameters of the filtering function. We note the following points regarding the entries in Table 1.
$\kappa$ $L$ $m$ $2m+1$ deg $\sf LB$} $\sf AI$} $\sf FAI$}
@@
    The number of variables of the filtering function $f_{2m+1}$ is $2m+1$ (and not $m$).
@@
    For each entry in the table, our calculation shows that $\beta\gg\gamma$. For example, for $\mathcal{S}(257,59)$, we obtained $\beta=2^{364.38}$ while $\gamma=2^{130.12}$. This is not surprising, since the fast algebraic attack is known to be much faster than the basic algebraic attack.
@@
    The feedback polynomial for the LFSR has to be a primitive polynomial of degree $L$. The above analysis of correlation attacks shows that for the obtained values of $L$ and $m$, it does not matter whether the polynomial is sparse, or whether it has a sparse multiple. So from an efficiency point of view, one may choose a low weight primitive polynomial. Concrete choices of primitive polynomials for the values of $L$ in Table 1 are shown in Table 2.
@@
Table 2: Concrete choices of primitive polynomials for the values of $L$ shown in Table 1.
@@
[…]
@@
$\mathcal{S}(L,m)$ $\sf posX$} and $\sf posY$}
$\mathcal{S}(161,37)$ d569a664f500763506c3
@@
$\mathcal{S}(161,37)$ 16 36
$\mathcal{S}(257,59)$ 26 57
$\mathcal{S}(331,71)$ 30 69
$\mathcal{S}(389,87)$ 39 86
$\mathcal{S}(449,101)$ 45 96
$\mathcal{S}(521,115)$ 51 112
@@
Given $\sf pos$} and the coefficient vector $\mathbf{c}$ of the primitive connection polynomial, it is required to determine the string $\mathbf{d}$ which provides the feedback positions for the initialisation round function $\sf IR$} (see (18)). The string $\mathbf{d}$ has to be chosen so as to satisfy the conditions given by (17). The weight of $\mathbf{d}$ is equal to $\mu=\lfloor\sqrt{2\kappa}\rfloor$. In Table 5 we provide concrete choices of $\mathbf{d}$ by listing the positions where $\mathbf{d}$ is 1.
$\mathcal{S}(L,m)$ $\mathbf{d}$
$\mathcal{S}(163,37)$ 30 42 55 67 84 90 104 114 130 138 150 162
$\mathcal{S}(257,59)$ 17 33 48 64 81 96 112 129 145 161 178 194 209 224 241 256
$\mathcal{S}(331,71)$ 25 43 61 78 98 115 134 151 168 186 206 224 242 259 277 294 312 330
$\mathcal{S}(389,87)$ 46 65 84 104 123 141 160 179 203 217 236 256 274 293 314 331 350 369 388
$\mathcal{S}(449,101)$ 29 49 70 94 112 133 155 175 197 218 239 260 282 303 322 344 364 386 408 431 448
$\mathcal{S}(521,115)$ 58 80 102 125 146 169 192 212 235 256 278 301 324 344 371 388 412 433 454 479 498 520
Table 5: Concrete choices of $\mathbf{d}$ corresponding to the various choices of $\sf pos$} obtained from $\sf posX$} and $\sf posY$} in Table 3, and the choices of the coefficient vector $\mathbf{c}$ in Table 2.
@@ 4.4 Differential Attacks
Choosing the filtering function to protect against certain known classes of attacks does not, however, protect against all possible attacks. There are attacks which can succeed even if the filtering function is properly chosen. One such attack on a previous version of the proposal was described in [6]. To describe the attack and the design modification to resist it we need to mention the aspects of the design which were different in the previous proposal. There are three such aspects.
@@
    In the definition of ${\sf MM}_{2m}$ in the earlier version we had chosen $\pi$ to be the identity permutation, whereas in the present version we choose $\pi$ to be the bit reversal permutation.
@@
In other words, due to the choice of $\pi$ as the identity permutation, the selection of tap positions as consecutive positions of the state, and the fact that two successive keystream bits are obtained from a single shift of the LFSR sequence, m1m-1 of the quadratic terms in the inner product cancelled out when $w_{t}$ and $w_{t+1}$ are XORed together. Further, the inputs to the two calls to $\sf Maj$} have an overlap of size m1m-1. These features were observed in [6] and exploited to mount a differential attack. The idea of the differential attack is to introduce a difference in the top most bit position of the IV. During the initialisation phase, this difference travels unchanged for $\kappa-m$ steps and then produces a difference in the input to the filtering function. The difference in the output of the filtering function is fed back into the leftmost bit of the LFSR. As the LFSR is further shifted, this difference then travels without any further modification creating a high probability truncated differential. Combined with the simplified form of $w_{t}\oplus w_{t+1}$, this leads to an efficient key recovery attack.
@@
Further, due to our design procedure for the tap positions corresponding to the variables $X_{1},\ldots,X_{n}$, the overlap of inputs in the calls to $\sf Maj$} in the XOR of any two keystream bits is at most $\nu$. From Table 4, we observe that the value of $\nu$ is considerably smaller than $m$. So the complexity of $\sf Maj$} is preserved in the XOR of any number of keystream bits.
Further, in the present version, during the initialisation phase we inject the output of the filtering function at multiple positions of the state, the number of positions being the weight of the string $\mathbf{d}$ (see (17)) which is about $\mu=\lfloor\sqrt{2\kappa}\rfloor$. Further, the positions of the 1’s in $\mathbf{d}$ are more or less equi-spaced so that the difference between two 1-positions is also about $\mu$. So any (controlled) difference in any bit position travels at most $\mu$ steps before it is further modified. In the full initialisation phase consisting of $2\kappa$ iterations, a difference will be updated about $\sqrt{2\kappa}$ times. This makes it difficult to control a difference through all the initialisation rounds.
@@ 5 Efficiency of Computing ${\sf MM}_{n}$
The computation of ${\sf MM}_{2m}$ requires computing ${\sf Maj}_{m}$ and an inner product of two $m$-bit strings. We discuss basic strategies for implementing these operations which provide estimates of the number of gates required.
@@
This operation requires $m$ AND gates and m1m-1 XOR gates.
Computation of ${\sf Maj}_{m}$.
The computation of ${\sf Maj}_{m}$ requires the computation of the weight of an $m$-bit string and the computation of a majority function on the weight. We first consider the computation of the majority function on the weight, and then we consider the computation of the weight of an $m$-bit string.
Majority function on the weight of an $m$-bit string.
We start with an example. From Table 1, one of the choices of $m$ is $m=37$. The weight of a 37-bit string is a 6-bit quantity, say $w_{5}w_{4}w_{3}w_{2}w_{1}w_{0}$. It is required to determine whether the value represented by this string is at least 19. This is computed by the Boolean formula $w_{5}\vee(w_{4}\wedge(w_{3}\vee w_{2}\vee(w_{1}\wedge w_{0})))$, requiring 3[OR]+2[AND] gates. In general, the weight of an $m$-bit string is an $\omega$-bit value, where $\omega=\lceil\log_{2}m\rceil$. To compute the threshold function, $\omega_{1}$ OR and $\omega_{2}$ AND gates are required for some values of $\omega_{1}$ and $\omega_{2}$ satisfying $\omega_{1}+\omega_{2}\leq\omega$. So the number of gates for computing the threshold function on the weight of an $m$-bit string requires a logarithmic (in $m$) number of gates.
@@
Using the above algorithm, it is easy to show that the computation of the weight of an $m$-bit string, when $m=2^{r}-1$ with $r\geq 1$, requires $2^{r}-r-1$ full adders. From this it easily follows that the number of full adders required to compute the weight of an $m$-bit string for arbitrary $m$ is $O(m)$. Since a full adder can be implemented using a constant number of NAND gates, the number of NAND gates required to compute the weight of an $m$-bit string is also $O(m)$. We record this fact in the following result.
@@
Concrete gate count estimates for $\mathcal{S}(L,m)$.
Given $L$ and $m$, the gate count estimate for $\mathcal{S}(L,m)$ primarily arises from the gate count estimates of the $L$-bit state and the filtering function $f_{2m+1}$. Beyond these two components, a small number of XOR gates are required to compute the next bit function $\sf nb$} of the LFSR (see Table 2) and $\mu=\lfloor\sqrt{2\kappa}\rfloor$ XOR gates are required by the initialisation round function $\sf IR$} (see (18) and the discussion following Proposition 5). From (6), the estimate for the gate count of $f_{2m+1}$ is $2$ plus an estimate of the gate count for computing $({\sf rev},{\sf Maj}_{m})\mbox{-}{\sf MM}_{2m}$. From the above discussion, an estimate of the gate count for ${\sf MM}_{2m}$ is $m+\lceil\log_{2}m\rceil$ AND/OR gates, and m1m-1 XOR gates, plus the number of gates required to compute the weight of an $m$-bit string.
@@
If we consider a flip-flop to be 12 units as done for Trivium [9], then the gate count estimate for $\mathcal{S}(163,37)$ is 2404.5. The gate estimate for Trivium obtained in [9] is 3488. Both Trivium and $\mathcal{S}(163,37)$ are targeted at the 80-bit security level, and $\mathcal{S}(163,37)$ is substantially smaller than Trivium. The lower size of $\mathcal{S}(163,37)$ is due to the smaller state size; $\mathcal{S}(163,37)$ uses a 163-bit state, while Trivium uses a 288-bit state.
The gate count estimate for Grain-128a obtained in [1] is 2145.5. Grain-128a is targeted at the 128-bit security level. Comparing to $\mathcal{S}(257,59)$, which is also targeted at the 128-bit security level, we find the gate count estimate of $\mathcal{S}(257,59)$ to be 2840.5. The state sizes of both Grain-128a and $\mathcal{S}(257,59)$ are almost the same. The greater size of $\mathcal{S}(257,59)$ is due to the greater gate size requirement of the filtering function $f_{2m+1}$ in comparison to the gate size requirement of the nonlinear components of Grain-128a. Even though $\mathcal{S}(257,59)$ is larger than Grain-128a, its size of about 2840.5 gates is small enough for $\mathcal{S}(257,59)$ to be considered as a small cipher. Finally we note that $\mathcal{S}(521,114)$ which is targeted at the 256-bit security level requires about 5693.0 gates. We know of no other 256-bit secure stream cipher which has such a small gate count.
We note that it is possible to improve security at the cost of a reasonable increase in gate count. For example, from Table 1, at the 128-bit security level, our proposal has $L=257$ and $m=59$, resulting in linear bias equal to $2^{-60}$ and the fast algebraic attack requiring more than $2^{130.12}$ operations. If we increase $m$ from 59 to 63, then the linear bias drops to $2^{-64}$, and the fast algebraic attack now requires more than $2^{135.83}$ operations. The gate estimate for $\mathcal{S}(257,63)$, i.e. the stream cipher with $L=257$ and $m=63$, is 2887.5 gates (the LFSR requires 2056 gates, and $f_{127}$ requires 776.5 gates). So the security increases by about 5 bits at the cost of an increase of only 47 gates in the circuit size.